# XOXO: Stealthy Cross-Origin Context Poisoning Attacks against AI Coding Assistants

### v1

@@ 7.3 RQ3: Code Reasoning
Clone Detection. Results in  Table IX indicate GCGS exceeds all existing approaches across models. On CodeBERT, GCGS achieves 72.27% success rate, surpassing the next best baseline RNNS by 29.40%. The warmed-up variant (GCGS+W) further rises to 80.97%. Similar gains are seen on GraphCodeBERT and CodeT5+, with GCGS reaching 64.02% and 65.34% respectively, and GCGS+W achieving 83.19% and 69.04%. While GCGS requires more queries than baselines like WIR (224-236 queries), the significantly higher success rates justify this. GCGS+W enhances efficiency with 35.34% fewer queries while boosting success rate, with GraphCodeBERT seeing the largest reduction from 1150 to 545 queries.
TABLE IX: Performance of GCGS and GCGS+W (warmed-up) attacks compared to SoTA baselines on CodeXGLUE Clone Detection. Results show mean ± std over 5 seeds. Each best ASR per model is bold.
CodeBERT GraphCodeBERT CodeT5+ Attack ASR # Queries ASR # Queries ASR # Queries ALERT 19.32 ±6.15 2125 ±161 21.02 ±2.89 2083 ±87 25.35 ±5.16 2008 ±117 MHM 26.10 ±8.98 1000 ±85 32.78 ±6.05 944 ±55 37.97 ±8.25 874 ±74 RNNS 42.87 ±4.27 1036 ±66 44.91 ±4.03 967 ±59 46.90 ±7.59 1045 ±109 WIR 24.76 ±6.48 236 ±15 30.41 ±6.01 224 ±7 31.78 ±6.36 224 ±12 GCGS 72.27 ±5.38 1032 ±106 64.02 ±6.70 1150 ±100 65.34 ±3.82 1078 ±42 GCGS+W 80.97 ±2.48 728 ±113 83.19 ±3.20 545 ±69 69.04 ±8.77 835 ±165
@@
In addition, GCGS generates more natural adversarial examples compared to other methods (see  Table X). On CodeBERT, GCGS modifies 4.13 identifiers across 15.70 positions with a CodeBLEU of 93.14, maintaining high similarity to original code. Warmed-up GCGS reduces modifications to 2.64 identifiers and 9.44 positions while raising CodeBLEU to 95.63, yielding both higher success rates and more natural adversarial examples.
@@
TABLE X: Naturalness of GCGS and GCGS+W (warmed-up) attacks compared to SoTA baselines on CodeXGLUE Clone Detection. Results show mean ± std over 5 seeds. Each best score per model is bold.
@@
CodeBERT GraphCodeBERT CodeT5+ Attack # Identifiers # Positions CodeBLEU # Identifiers # Positions CodeBLEU # Identifiers # Positions CodeBLEU ALERT 4.46 ±0.87 18.56 ±3.53 84.34 ±2.09 4.25 ±0.60 18.37 ±2.44 83.13 ±1.78 3.58 ±0.40 15.06 ±3.69 86.35 ±1.94 MHM 5.71 ±0.23 24.39 ±2.01 84.04 ±2.07 5.84 ±0.30 24.64 ±1.28 84.71 ±0.63 4.89 ±0.25 19.97 ±1.29 86.72 ±0.87 RNNS 5.87 ±1.01 25.98 ±6.53 92.37 ±1.17 5.34 ±0.80 23.12 ±2.56 92.76 ±0.94 4.04 ±1.10 19.73 ±5.71 93.37 ±1.19 WIR 5.03 ±0.69 21.70 ±2.60 87.40 ±1.19 4.82 ±0.23 21.31 ±1.43 87.61 ±0.88 3.96 ±0.32 16.47 ±1.47 89.46 ±0.85 GCGS 4.13 ±0.35 15.70 ±1.43 93.14 ±0.53 3.81 ±0.40 14.90 ±1.65 93.76 ±0.35 2.79 ±0.24 11.29 ±1.52 94.56 ±0.63 GCGS+W 2.64 ±0.23 9.44 ±1.26 95.63 ±0.72 2.05 ±0.22 7.75 ±0.71 96.38 ±0.43 1.98 ±0.29 7.20 ±1.96 96.66 ±0.75
{longfbox}
Result 3: Warmed-up GCGS outperforms SoTA adversarial attacks leveraging semantics-preserving code transformations on security-relevant code reasoning tasks by up to 38.28 percentage points.
7.4 RQ4: Cross-Model Warm-up
While warming up GCGS improves both performance and naturalness, it requires an initial query investment that must be amortized over multiple attacks. We therefore investigate whether this cost can be eliminated by learning from a surrogate model rather than the target model itself. For each model in the Defect Detection dataset, we evaluate warm-up on the other two models as surrogates. Table XI and Table XII present our findings.

### v2

@@ 6.3 Code Reasoning
Clone Detection Results. GCGS exceeds all existing approaches across all models (Table 2). On CodeBERT, GCGS achieves 72.27% ASR, surpassing the next best baseline RNNS by 29.40 percentage points. The warmed-up variant (GCGS+W) further increases ASR to 80.97%. While GCGS requires more queries than baselines like WIR-Random (224-236 queries), the significantly higher ASR justifies this. GCGS+W make 52.61% fewer queries compared to GCGS while boosting ASR.
7 Related Work
Adversarial Robustness Evaluation. Frameworks performing adversarial robustness evaluation sample perturbations from an adaptive distribution using feedback from the targeted model.
Natural Language Transformations. Recent work has explored adversarial robustness using natural language prompt transformations [33, 64], assuming a threat model where attackers control IDE extensions to inject malicious prompt edits. In contrast, XOXO attack assumes a more practical threat model relying only on standard code-based transformations introduced through code changes to the target project, without requiring malicious extensions or prompt manipulation.
Whitebox Approaches. Many works rely on white-box model access, using feedback such as gradient information to guide attacks [67, 70, 13, 55, 49]. This requirement poses challenges for attacking large frontier models where attackers have only remote network access. GCGS uses a black-box approach, circumventing this requirement, effectively targeting frontier models.
Blackbox Approaches. A substantial body of black-box approaches has focused solely on code reasoning tasks [66, 69, 68, 43, 20, 58, 73, 39], relying on heuristics like model confidence in classification tasks such as defect and clone detection. In contrast, code generation requires reasoning over multiple tokens, rendering these approaches computationally infeasible. GCGS overcomes this with a lightweight method, effectively supporting code reasoning and generation tasks.
8 Limitations
Our work exposes significant vulnerabilities in AI-assisted software development, but the scope of our attack remains underexplored. We use identifier replacement as a semantics-preserving transformation, but the effectiveness of GCGS with other transformations is unclear. Additionally, GCGS’s greedy composition strategy relies on a monotonic error increase in the Cayley Graph, which may not apply to all model architectures. While the attack is difficult to detect during preprocessing due to benign modifications, we have not addressed post-processing guardrails that might filter vulnerabilities through token-level filtering. We outline potential defenses and their limitations in § B.
9 Conclusion
This paper introduces Cross-Origin Context Poisoning (XOXO), a novel attack that exploited automatic context inclusion in AI coding assistants, exposing a critical vulnerability in their architectures. Greedy Cayley Graph Search (GCGS) effectively applied semantics-preserving transformations that severely degraded the performance of leading generative and fine-tuned code LLMs, achieving an average 83.67% ASR on code generation and 84.51% on reasoning tasks. These findings underscore the need for robust defenses against context manipulation attacks.
Broader Impact Statement
Our findings highlight the urgent need for developing defenses against Cross-Origin Context Poisoning attacks. Furthermore, we hope our work raises user awareness of the potential risks AI coding assistants pose, encouraging users to critically evaluate suggested code to prevent introducing vulnerabilities into their software projects.
Appendix A Implementation
Transformations. Although the Cayley Graph structure accommodates any semantics-preserving transformations (including non-commutative ones), for attack implementation we focus on identifier replacements, specifically function, parameter, variable, and class-member names. This is because identifier replacements offer a larger search space compared to other transformations like control flow modifications, while enabling precise atomic control over the magnitude of code changes. We leverage tree-sitter [10] to parse code snippets and extract identifier positions. To maintain natural and realistic transformations, we employ different identifier sourcing strategies for each task.
For defect and clone detection tasks, we seed identifiers from their respective training sets to avoid out-of-distribution effects in fine-tuned models. For smaller Python datasets (HumanEval+, MBPP+, and CWEval/Python), we extract identifiers from CodeSearchNet/Python [31] to ensure sufficient variety. HumanEval+ and MBPP+ tasks additionally incorporate Python input-output assertions in docstrings (e.g., $>$ string_xor(’010’, ’110’) ’100’ or assert is_not_prime(2) == False), we maintain consistency by replacing function names in both the code and assertions as done by previous implementations [60, 28]. This consistency is crucial as the assertions are part of the model’s input, and any naming discrepancies would test the model’s ability to handle inconsistent references rather than its code understanding. When composing transformations (as illustrated in Figure 3), we iterate through identifier-replacement pairs ordered by increasing model confidence (based on the stored $g$-$\alpha$ map). For classification tasks, we measure the model’s confidence as the probability of predicting the correct class. For generation tasks, we measure the model’s confidence as the sum of the generated tokens’ log probabilities. At each iteration we select the highest-confidence pair where neither the identifier nor its replacement appears in previous steps. This process continues until we either discover a breaking transformation or exhaust the maximum number of queries to the model.
Warm Up. To highlight the practicality of attack warm-up, we use a small (less than 5% of the dataset) sample of the model’s training and validation datasets for reasoning tasks, illustrating that an attacker requires minimal access to in-distribution examples for effective results. This set ($C^{W}$) is kept disjoint from the model’s fine-tuning set to ensure fair evaluation.
For code completion tasks, where dataset sizes are smaller, we implement cross-dataset warm-up (e.g., using MBPP+ to initialize attacks on HumanEval+ and vice versa). The warm-up process begins with randomly sampling replacements for each identifier in the training set code snippets ($C^{T}$) and tracking the average drop in model confidence for each replacement across the complete $C^{T}$. Based on the top performing replacements from $C^{T}$, an attack is executed on $C^{V}$ for getting each replacement’s validation performance score. Using this score, we select top-$k$ highest-scoring transformations as warm-up set for the actual attack. We also experimented with alternative sampling methods, including distribution biasing and softmax-based sampling, but found that the straightforward top-$k$ selection strategy provided the best results.
Appendix B Defenses
Static Code Analysis. Static code auditing tools can serve as a defense measure either during code generation or as a post-generation phase. However, these tools currently face critical limitations [35, 34, 46, 38, 7, 8] that undermine their ability to be an effective defense strategy. First, due to the stringent latency requirement of code generation, existing tools require lightweight analysis (i.e., small ML models or regex/pattern matching) that sacrifices accuracy for low latency [7, 8, 23]. Second, post-generation tools scanning entire repositories often produce excessive false positives [35, 34, 46, 38]. Third, both approaches struggle with logical vulnerabilities that require manually provided, precise, application-specific specifications. Our CWEval evaluation shows GCGS can trigger logical vulnerabilities (see § I for details), which are extremely hard for code auditing tools to detect.
Adversarial Fine-tuning. Although successful in other domains, adversarial fine-tuning has been ineffective against our attacks. Our experiments show that even after fine-tuning with adversarial examples, models remained vulnerable, with ASR above 87% across all tested models (Table 11). In some cases, such as with CodeBERT and GraphCodeBERT, attack effectiveness even increased after fine-tuning. We speculate that this might be an effect of the smaller sizes of these models.
Appendix C Additional Evaluation
C.1 Experimental Details
Model Confidence Guidance. For code generation, while token-level probabilities may be accessible, they do not directly reflect the model’s confidence or output quality, making them unreliable signals for guiding attacks. Furthermore, SoTA adversarial methods designed for code reasoning tasks become computationally infeasible for generative tasks, as each evaluation requires generating a full sequence rather than a single classification decision. Therefore, for generative tasks, we cannot compare GCGS to the same methods we compared against in code reasoning.
We assess all open-source models with and without perplexity feedback. For Claude 3.5 Sonnet v2, we evaluate GCGS without perplexity due to their API restrictions on accessing token-level log probabilities. In the case of GPT 4.1, we solely evaluate GCGS+P with perplexity due to budget constraints and fewer queries made by GCGS when using perplexity as feedback. Due to cost considerations with closed-source models and limited compute resources for open-source models, we exclude GCGS with warm-up (GCGS+W) in the generation setting.
Machine Details. We conducted model fine-tuning using consumer hardware: a 20-core processor with 64GB RAM and dual NVIDIA RTX 3090 GPUs, running Ubuntu 22.04 and CUDA 12.1 (machine A). For in-context code generation and vulnerability injection tasks, we utilized AWS EC2 p5e.48xlarge instance equipped with 192 cores, 2048GB RAM, and eight NVIDIA H200 GPUs (one GPU per attack) on Ubuntu 22.04 with CUDA 12.4 (machine B). For comparative evaluations against SoTA attacks, model transferability and adversarial fine-tuning experiments, we utilized GCP g2-standard-96 instances equipped with 96 cores, 384GB RAM, and eight NVIDIA L4 GPUs (one GPU per attack) on Debian 11 with CUDA 12.1 (machine C). To serve LLMs, we use either transformers 4.42.4 [63] or vllm 0.6.3.post1 [37]. We access GPT 4.1 through OpenRouter and Claude 3.5 Sonnet v2 through GCP Vertex AI API.
Execution Time. For the final evaluation runs, we spent 17.22 GPU-days on model fine-tuning on machine A, 20.65 GPU-days on in-context code generation and vulnerability injection tasks on machine B, and 14.21 GPU-days on code reasoning attacks on machine C. We have spent about 1.5 days running experiments on Claude 3.5 Sonnet v2 through GCP Vertex AI API and another 1.5 days running experiments on GPT 4.1 through OpenRouter. We estimate that total usage, including reruns and development, might be 2-3 times higher than our evaluation runs.
C.2 Model and Dataset Licenses.
We include the licenses for models in Table 4 and datasets in Table 4.
Table 3: License information for the evaluated models.
Model License Claude 3.5 Sonnet v2 (2024/10/22) Proprietary GPT 4.1 (2025/04/14) Proprietary Codestral 22B v0.1 mnlp-1.0 DeepSeek Coder 6.7B Instruct Deepseek License DeepSeek Coder 33B Instruct Deepseek License Llama 3.1 8B Instruct Llama3.1 Community License Qwen 2.5 Coder 7B Instruct Apache-2.0 Qwen 2.5 Coder 32B Instruct Apache-2.0 CodeBERT MIT GraphCodeBERT MIT CodeT5+ 110M BSD-3
Table 4: License information for the datasets employed.
Dataset License HumanEval+ Apache-2.0 MBPP+ Apache-2.0 CodeXGlue Creative Commons v1.0
C.3 Baseline Model Performance
We evaluated baseline performance of models on the code generation, vulnerability injection (both in Table 5), and reasoning tasks ( Table 6).
While we considered evaluating generative coding models in a zero-shot chat setting, our experiments (shown in the upper part of Table 6) revealed they did not perform well (in fact close to random guessing, i.e., 50% accuracy) on both binary classification tasks, even with extensive prompt engineering. This poor baseline performance, which aligns with existing findings on LLMs’ limitations in vulnerability detection [17], led us to focus our evaluation on fine-tuned LLMs that show meaningful accuracy on these tasks.
Table 5: pass@1 performance of tested SoTA LLMs on code generation (HumanEval+ and MBPP+) and vulnerability injection (CWEval/Python).
Model HumanEval+ MBPP+ CWEval/Python Claude 3.5 Sonnet v2 (2024/10/22) 70.73 67.29 40.00 GPT 4.1 (2025/04/14) 80.49 77.13 48.00 Codestral 22B 75.00 57.71 48.00 DeepSeek Coder 6.7B 67.07 46.81 36.00 DeepSeek Coder 33B 70.73 65.16 48.00 Llama 3.1 8B 50.61 43.62 40.00 Qwen 2.5 Coder 7B 79.88 73.94 48.00 Qwen 2.5 Coder 32B 87.20 75.53 52.00
Table 6: Performance comparison (accuracy %) of zero-shot generative models against fine-tuned classifier models.
Zero-shot Generation Defect Detection Clone Detection Codestral 22B 55.73 53.75 DeepSeek Coder 6.7B 47.49 53.75 DeepSeek Coder 33B 54.75 53.50 Llama 3.1 8B 44.40 50.00 Qwen 2.5 Coder 7B 55.82 50.50 Qwen 2.5 Coder 32B 56.14 50.75 Fine-tuned Classifiers Defect Detection Clone Detection CodeBERT 62.03 ±0.88 90.05 ±1.33 GraphCodeBERT 62.95 ±0.62 97.30 ±0.19 CodeT5+ 110M 61.74 ±1.07 84.95 ±2.06
C.4 Attack Naturalness
We evaluate the quality and naturalness of adversarial examples using three metrics widely adopted in prior work [66, 71, 20]. (i) CodeBLEU [51] measures code similarity by combining BLEU score with syntax tree and data flow matching, ranging from 0 (completely distinct) to 100 (identical). Higher scores indicate adversarial code that better preserves the original code’s structure and functionality. (ii) and (iii) Identifier and Position Metrics (# Identifiers, # Positions) count the number of replaced identifiers and their occurrences in the code. For instance, changing one variable used multiple times affects several positions. Lower numbers indicate more natural modifications that are harder to detect through static analysis or code review.
Table 7: Naturalness of GCGS and GCGS+P (perplexity-guided) attacks on code generation using HumanEval+ and MBPP+. Results show mean ± std over 5 seeds for open-source models and single runs for closed-source models (limited 5-run analysis in § H).
HumanEval+ MBPP+ Model Attack # Identifiers # Positions CodeBLEU # Identifiers # Positions CodeBLEU Claude 3.5 Sonnet v2 GCGS 1.00 2.86 98.50 1.00 1.82 98.53 GPT 4.1 GCGS+P 2.30 6.11 97.52 3.48 8.13 94.72 Codestral 22B GCGS 1.00 ±0.00 1.60 ±0.16 98.83 ±0.09 1.00 ±0.00 1.29 ±0.04 99.01 ±0.03 GCGS+P 2.01 ±0.18 4.38 ±0.51 97.87 ±0.20 1.17 ±0.07 1.81 ±0.23 98.68 ±0.15 DeepSeek Coder 6.7B GCGS 1.00 ±0.00 1.84 ±0.08 98.08 ±0.06 1.00 ±0.00 1.74 ±0.08 98.40 ±0.09 GCGS+P 2.27 ±0.19 5.45 ±0.62 96.98 ±0.45 1.14 ±0.06 2.13 ±0.21 98.16 ±0.11 DeepSeek Coder 33B GCGS 1.00 ±0.00 2.01 ±0.31 98.81 ±0.15 1.00 ±0.00 1.42 ±0.08 98.89 ±0.05 GCGS+P 2.60 ±0.31 6.54 ±0.73 97.22 ±0.26 1.27 ±0.07 2.24 ±0.23 98.39 ±0.14 Llama 3.1 8B GCGS 1.00 ±0.00 1.69 ±0.17 98.80 ±0.11 1.00 ±0.00 1.66 ±0.10 98.59 ±0.08 GCGS+P 1.84 ±0.16 4.37 ±0.56 97.91 ±0.21 1.11 ±0.07 1.90 ±0.17 98.43 ±0.12 Qwen 2.5 Coder 7B GCGS 1.00 ±0.00 1.24 ±0.06 99.06 ±0.06 1.00 ±0.00 1.26 ±0.03 99.06 ±0.03 GCGS+P 1.78 ±0.24 3.54 ±0.79 98.27 ±0.24 1.63 ±0.11 3.19 ±0.19 97.93 ±0.09 Qwen 2.5 Coder 32B GCGS 1.00 ±0.00 1.48 ±0.15 99.04 ±0.07 1.00 ±0.00 1.32 ±0.04 98.96 ±0.05 GCGS+P 2.39 ±0.21 5.16 ±0.71 97.80 ±0.25 1.45 ±0.09 2.51 ±0.37 98.22 ±0.25
Code Generation and Vulnerability Injection. In Table 7 and Table 8, we see the adversarial examples maintain high naturalness across all models, as evidenced by CodeBLEU scores consistently above 96. The base GCGS achieves slightly higher CodeBLEU due to the limited modification scope. In contrast, perplexity-guided GCGS (GCGS+P) makes more extensive but still natural modifications, affecting more identifiers and positions while maintaining comparable CodeBLEU scores. This suggests that GCGS+P finds a better balance between attack effectiveness and naturalness.
Defect Detection. As shown in Table 9, GCGS outperforms baselines in code naturalness, averaging only 1.84 identifier changes and 9.65 position modifications. Likewise, its average CodeBLEU score of 92.94 exceeds WIR-Random’s 86.05. With warm-up, GCGS+W further improves, requiring 1.05 identifier and 2.61 position changes when attacking CodeT5+.
@@ C.4 Attack Naturalness
Clone Detection. GCGS generates more natural adversarial examples compared to other methods (see  Table 10). On CodeBERT, GCGS modifies 4.13 identifiers across 15.70 positions with a CodeBLEU of 93.14, maintaining high similarity to original code. Warmed-up GCGS reduces modifications to 2.64 identifiers and 9.44 positions while raising CodeBLEU to 95.63, yielding both higher success rates and more natural adversarial examples.
@@
Table 8: Naturalness of GCGS and GCGS+P (perplexity-guided) attacks on CWEval/Python. Results show mean ± std over 5 seeds. Each best score per model is bold.
Model Attack # Identifiers # Positions CodeBLEU Claude 3.5 Sonnet v2 GCGS 1.00 1.75 99.37 GPT 4.1 GCGS+P 1.00 1.17 99.66 Codestral 22B GCGS 1.00 ±0.00 1.46 ±0.25 99.30 ±0.03 GCGS+P 2.00 ±0.34 3.60 ±0.53 98.69 ±0.18 DeepSeek Coder 6.7B GCGS 1.00 ±0.00 1.07 ±0.10 99.60 ±0.04 GCGS+P 1.50 ±0.65 1.95 ±1.42 99.37 ±0.41 DeepSeek Coder 33B GCGS 1.00 ±0.00 1.33 ±0.19 99.48 ±0.10 GCGS+P 1.38 ±0.55 2.16 ±1.30 99.28 ±0.34 Llama 3.1 8B GCGS 1.00 ±0.00 1.75 ±0.30 99.48 ±0.08 GCGS+P 2.68 ±1.02 4.82 ±2.08 98.83 ±0.43 Qwen 2.5 Coder 7B GCGS 1.00 ±0.00 1.80 ±0.23 99.37 ±0.05 GCGS+P 2.85 ±1.37 6.27 ±3.16 98.44 ±0.67 Qwen 2.5 Coder 32B GCGS 1.00 ±0.00 1.18 ±0.29 99.50 ±0.14 GCGS+P 2.90 ±2.52 4.67 ±5.07 98.66 ±1.13
Table 9: Naturalness of GCGS and GCGS+W (warmed-up) attacks compared to SoTA baselines on CodeXGLUE Defect Detection. Results show mean ± std over 5 seeds. Each best score per model is bold.
@@
CodeBERT GraphCodeBERT CodeT5+ Attack # Identifiers # Positions CodeBLEU # Identifiers # Positions CodeBLEU # Identifiers # Positions CodeBLEU ALERT 3.01 ±0.23 25.42 ±1.85 81.59 ±1.72 2.62 ±0.22 20.04 ±2.34 84.35 ±0.87 2.95 ±0.25 23.68 ±1.22 82.91 ±0.56 MHM 2.74 ±0.30 20.54 ±2.79 84.67 ±1.19 2.59 ±0.21 17.88 ±2.08 86.05 ±0.78 2.75 ±0.16 19.45 ±1.58 84.18 ±0.75 RNNS 3.92 ±0.59 32.45 ±6.23 86.85 ±1.10 2.60 ±0.49 22.43 ±4.66 88.01 ±0.88 2.76 ±0.29 23.53 ±3.59 87.74 ±0.94 WIR-Random 2.64 ±0.17 21.99 ±2.03 85.05 ±0.99 2.22 ±0.26 17.22 ±2.55 86.91 ±0.72 2.40 ±0.20 18.39 ±1.28 86.18 ±0.73 GCGS 2.00 ±0.26 9.96 ±2.33 92.83 ±1.28 1.94 ±0.24 11.51 ±2.55 91.61 ±1.29 1.57 ±0.06 7.48 ±0.53 94.37 ±0.38 GCGS+W 1.49 ±0.18 5.94 ±1.38 95.62 ±0.94 1.45 ±0.26 7.83 ±2.85 94.13 ±2.00 1.05 ±0.03 2.61 ±0.90 97.93 ±0.77
Table 10: Naturalness of GCGS and GCGS+W (warmed-up) attacks compared to SoTA baselines on CodeXGLUE Clone Detection. Results show mean ± std over 5 seeds. Each best score per model is bold.
CodeBERT GraphCodeBERT CodeT5+ Attack # Identifiers # Positions CodeBLEU # Identifiers # Positions CodeBLEU # Identifiers # Positions CodeBLEU ALERT 4.46 ±0.87 18.56 ±3.53 84.34 ±2.09 4.25 ±0.60 18.37 ±2.44 83.13 ±1.78 3.58 ±0.40 15.06 ±3.69 86.35 ±1.94 MHM 5.71 ±0.23 24.39 ±2.01 84.04 ±2.07 5.84 ±0.30 24.64 ±1.28 84.71 ±0.63 4.89 ±0.25 19.97 ±1.29 86.72 ±0.87 RNNS 5.87 ±1.01 25.98 ±6.53 92.37 ±1.17 5.34 ±0.80 23.12 ±2.56 92.76 ±0.94 4.04 ±1.10 19.73 ±5.71 93.37 ±1.19 WIR-Random 5.03 ±0.69 21.70 ±2.60 87.40 ±1.19 4.82 ±0.23 21.31 ±1.43 87.61 ±0.88 3.96 ±0.32 16.47 ±1.47 89.46 ±0.85 GCGS 4.13 ±0.35 15.70 ±1.43 93.14 ±0.53 3.81 ±0.40 14.90 ±1.65 93.76 ±0.35 2.79 ±0.24 11.29 ±1.52 94.56 ±0.63 GCGS+W 2.64 ±0.23 9.44 ±1.26 95.63 ±0.72 2.05 ±0.22 7.75 ±0.71 96.38 ±0.43 1.98 ±0.29 7.20 ±1.96 96.66 ±0.75
C.5 Adversarial Fine-tuning
We investigate whether adversarial fine-tuning can effectively defend against GCGS attacks. Following established approaches in adversarial attack literature [66, 29], we augment the target models’ training sets with adversarial examples. For each model (CodeBERT, GraphCodeBERT, and CodeT5+), we first generate adversarial examples from the Defect Detection training set using GCGS as follows: for each training set example, we either generate a single adversarial example or, if the attack on a particular example was unsuccessful, we use the example where the target model was the least confident about the correct class. We then create an adversarially-augmented training set by combining and shuffling the original training data with these adversarial examples. After fine-tuning each model on their respective augmented training sets, we evaluate this defense by running GCGS against the fine-tuned models.
Table 11: GCGS results on GCGS-adversarially fine-tuned models.
Model ASR # Queries # Identifiers # Positions CodeBLEU CodeBERT 99.35 57 1.32 4.61 96.32 GraphCodeBERT 99.93 30 1.18 5.38 95.87 CodeT5+ 87.42 444 2.28 12.96 90.81
Table 11 presents our findings. Adversarial fine-tuning proves ineffective against GCGS across all tested models. For CodeBERT and GraphCodeBERT, the attack’s effectiveness and efficiency actually appear to increase after fine-tuning, though this may be attributed to experimental variance. Even in the best case, with CodeT5+, adversarial fine-tuning only reduces attack effectiveness by 10.34 percentage points while decreasing efficiency by a factor of 2.51 – far from preventing the attack. These results suggest that the impact of adversarial fine-tuning heavily depends on the underlying model architecture, and even in optimal conditions, fails to provide meaningful protection against GCGS attacks.
C.6 Cross-Model Warm-up